Tech Staff Meeting Notes March 9, 2010

The top priority is always repair of down stations.
NM problem stations
ET problem stations

The task list updated February 6, 2009 needs to be updated again. Note that everyone is still needed to work on essentially all tasks. There is however, a single person who is primarily responsible for each one. Mitch of course, is ultimately responsible for everything.


  1. DBB/lnxt rack repair
  2. Repair down stations and routine maintenance
  3. GNAR repair
  4. FPAL
  5. NM broadband ARRA upgrade prep

News & Highlights

The Research Equipment Tech I position is filled. Please welcome David Steiner.

Friday March 12 is a University staff Holiday. Everyone else is on spring break this week.

Chris is arranging annual safety training for himself, David, and Jim. Unfortunately none of the staff were able to attend so Chris will arrange another session.


Short Period & Broadbands

Holly's NM list:
High priority stations: RDGT, WYBT, MIST, MARM, GOBM and TNMT.

Data taken from:

Muted Amplitude

Isis Oxidation

Gain Channel problems

Unusable unless it's a large event

Jim noted that single component problems are non-trivial to fix. He asked if some stations could be decommissioned and the hardware resinstalled at bad stations. Mitch asked whether new hardware could be built and used to replace aging systems--the design is sound but some hardware is nearly 15 years old.

NODE Tasks

Richardson, LNXT, and SFTN need new batteries.


Need alternate comm path. Telecom says dsl isn't available (though we once had ISDN).

Mitch met with Bob Woodward and Bob Busby of IRIS about EarthScope. Bob Busby said he might be able to loan us some WildBlue hardware to test. Assuming we can get sufficient bandwidth without violating shared use maxima, this could be a viable option if dsl is unavailable.


Still need to tie generator status alarms into lnxtworm

Now that we have a phone line for the dsl, we can replace the adam6060 with the simpler style phone reset.

The dsl on dual ethernets on lnxtworm causes conflicts and instability in the DBB. Steve attempted configuring a prioritized path without success. A second machine was installed which, combined with a bad switch, caused a cascade of power failures. Repairs should be complete by Wednesday or Thursday. Still need to test that the conflicts don't reappear when the 10.156 side is enabled on lnxtdsl.


Need alternate comm path. Telecom says dsl isn't available (though we once had ISDN).

AZTN: Upgrade PC and replace XLAK clock


Need to gather up a contact list (e.g. everyone's home phone, power companies, etc) and put it on webworm (password protected). Jim has a draft of this that so far has all the node contacts. Mitch suggested that the umdrive seisnet space might be a better place than webworm for it.

The DBB firewall installed and operating. Only specific ip's are allowed through. You need to use the route command to manually add a route. Evertying behind the firewall is the same D ip number but with 10.156.1 instead of 141.225.156. Remaining items include:


For some stations (most frequently SEAR) data feeds get dropped. Scream sits there "connecting" but never connects. The mss100 is fine. A disconnect then connect on scream fixes the problem (or rather fixes the symptom--the problem is unknown). Using reboot_mss_ew didn't help. We still have problems at SEAR even with a new mss100. While we still want to upgrade everything to year 3 hardware, Mitch will seek permission to stream at 50 sps and trigger at 200. Jim will verify that we can actually do that. No response from ANSS management on request to stream at 50 and trigger at 200.

NAIT: Couldn't get the NWS cable modem to cooperate at NAIT (firewall?). Dean was going to ask Comcast about it and maybe get a new modem. Steve will follow up.

CVVA is upgraded but there remains an unresolved issue with the firewall Steve is working with Merritt Reagan NGC to get them resolved.

Al is making a practice run on the strongmo trigger download for data from the 3.4 near Dyersburg on Nov 6. Mitch sent these instructions are archving triggered strongmotion data:

Recorders 1, 6, and 7 are repaired, configured, and awaiting installation. Al reports that of the remaining four, only two are associating.

Resets are not working well and tones are degraded. Likely cause is telemetry between the reset unit on the AZTN roof and the DAS's on the bridge. If its ice, it will self-correct. Otherwise we'll check it when we get time and a lane closure to install the 3 repaired units.


DSL is running at Boswell and we can see the PC there (bowlworm). The VLAN problem is resolved and we can now see the dataloggers at HDAR and HDBT. I guess we have to fix the problems now and get things online.

Bad CMG5 E/W.

Mud Island vertical array is operational. Not really, functional is a better word until we can get it online to qc/vet data (see HDAR).

Replace two bad mss100's.

Bad 100' accelerometer.


Holly's ET list:
Data taken from:


Unusable - see attached data snippets

Also no alarms present in BRBC data.
Also no alarms present in BRBC data.

Other stations Down

FPAL is on deck for the installation of a RT130, T120, and episensor. Jim and David will get started on the panel after lnxt repairs are complete. Jim asked if it should have a router like the AG stations since we have no radios and hence, only need a single IP. Mitch suggested that we should since we don't know how robust the RT130 interface will be "in the wild" on the internet. Mitch sent an email to the netops list asking for advice.

Many of the NC station problems are probably because of insufficient power. John will find out what's needed to increase solar and battery to help make it through the winter storms. The ET stations have the same issues but problems are not as evident because they are, in general, at much lower elevation (less snow).


BHT. Chris reports that it is time to seriously pursue another site as access is even difficult by 4-wheeler. David reports moving it will take all the fun out it, what's the worst that could happen? John said that folks at Cumberland Trail State Park said we can put the station at another spot on Black Mtn. which is just a couple of miles NE of the present location. It is accessible by a paved road. We can pursue that or I can look for something that will put it further NE to get it more equidistant between DYTN and WMTN. Right now it is awfully close to DYTN and BCRT. Mitch likes the idea of splitting the difference between DYTN and WMTN. John will dig out his old notes and starting looking around.


GFM The NC state park service is purchasing GFM. We already have a couple permissions from them and John initiated the paperwork for the GFM permission.

TRYN: the forest service wants to remove the tower. There will be a new 180' tower and we have permission to be on it. We also have permission to be in the building (rack mount only). Can we convert this station to AC?

NCCU: Done! Thanks John and Steve.


James has a very nice SOH website at

MAIR is down. Needs a site visit.

All stations except MACC got a routine service by Chris.

STLE, NMKM, and RLAP all need a charger upgrade. They are currently working (and have been for a long time) so this is not necessarily a top priority.

Both receivers were replaced at HCES. Coax for Mattioli's receiver is trashed--awaiting new cable from them.

AG Network

The AG Network is complete. Data feed to NEIC is complete. Christy is working on dataless afterwhich Mitch will start archiving at IRIS. Jim has a few loose ends he'd like to finish.

ARRA Upgrades

We are approved for the following in approximately the same order as we hope to complete them:

  1. Upgrade 11 NM broadband stations with new 6-c DAS, new sensor, and accelerometer.
  2. Upgrade 2 ET broadband stations with new 6-c DAS, new sensor, and accelerometer (CPCT and SWET).
  3. Upgrade 3 ET shortperiod analog stations with new 6-c das, digital telemetry, and accelerometer (BCRT, GMG, and GTTN).

ASL is sending the 13 Reftek's and 13 Trilliums. ASL ordered cables. No word on accelerometers nor additional 3 refteks.

Note that using other funds, we're also upgrading 3 additional ET stations to digital with episensor and FPAL to the same hardware as the other broadbands.

Mitch hung a chart in the war room to track our progress (we're required to make quarterly progress reports and to meet the proposed schedule).

We need to be very careful to keep AG, ARRA, and CERI hardware separate and be certain that hardware purchased by each project goes only to that project.

All replaced DM24's and CMG40T's are to go to south campus and taken off inventory. This hardware is not to be redeployed anywhere in the CERI seismic network nor FDL.


Jer-Mings vault construction is in progress.

Two CMG6TD's are in need of depot level repair. Mitch will see about availability of funds to send them to the guralp site in Menlo.

Computer Room Power

Need to review power loads and line balance after adding the new AQMS servers. Review needs to consider addition of new landmark servers, mac server, and seisnet deep waveform buffers. The Earth Science Library (ESL) in 3892 is moving to the main branch and has a liebert. We should consider carving out rack space in that room (though backup power is an issue).


The neic has a new metadata website at John is populating it with info from our stations.

Mitch is working on the mountain of paperwork required to contract help from Dave Ketchum to get initial dataless SEED from INV. It will include a visit to Memphis for training. John will be our INV czar.


Our cell phones have extremely poor performance in outlying areas. They should be replaced.

Tom Brackman at NKU has a shiny new CMG6TD that he wants to install and get integrated with ANSS. He'll run scream and Steve will help him get a minimal earthworm going to export data (only need startstop, statmgr, scream2ew, and export_scnl).

What to do with the Steiner brain dump DVD's?

Mitch made a seisnet group on umdrive. Go to You will need your uuid and password. There are several things on there including John's database. We can put contact info there too when its available. Only people in the seisnet group have access to this area. If you download something for update, don't forget to upload it again so that we all have the same version of the files.

There is now a ceri_seis email group that includes all the seismic networks staff.

Rotational sensor is feeding an earthworm but still has IRIG crosstalk problem to fix. Its fine on an Oscope hooked to the output of the mux and looks fine in LabView. Its bad when viewed with wave_viewer. Need to see what the sac files look like to make sure it isn't just a wave_viewer feature.

Chris continues to hound the supply to try to get Parker's old Suburban onto govdeals.

The pink palace display now includes a heli connection for MPHB. Need to retrieve any remaining hardare (e.g. antenna and reciever on roof).


3890 Power Chris is going to make a diagram for 3890. Good luck son.

3876 Power Chris is going to make a diagram for 3876. Good luck son.

Should we have a trouble ticket system?